If Z1 =5i-2j
Z2= 3i+3j
Z3=4i-j,
Show that Z1,Z2 and Z3 are associative, commutative and distributive.
1
Expert's answer
2020-12-14T07:25:46-0500

For associative property, elements of set must satisfy following condition :

"Z_i+(Z_j+Z_k)=(Z_i+Z_j)+Z_k" , where "Z_i,Z_j \& Z_k" are the different elements of a given set.

let's check..

"Z_1+(Z_2+Z_3)\stackrel{?}{=}(Z_1+Z_2)+Z_3"

LHS:

"LHS:5\hat i-2\hat j+(3\hat i+3\hat j+4\hat i-\hat j)=12\hat i\\
RHS:(5\hat i-2\hat j+3\hat i+3\hat j)+4\hat i-\hat j=12\hat i"

Hence, Z2,Z2 and Z3 are associative under addition operation.

For commutative property, we must have ...

"Z_i+Z_j=Z_j+Z_i" for all i and j.

let's check...

"Z_1+Z_2\stackrel{?}{=}Z_2+Z_1"

"LHS: (5\hat i-2\hat j)+(3\hat i + 3\hat j)= 8\hat i+\hat j\\
RHS: (3\hat i + 3\hat j)+(5\hat i-2\hat j)= 8\hat i+\hat j"

similarly we can show that...

"Z_1+Z_3=Z_3+Z_1=9\hat i-3\hat j\\
Z_3+Z_2=Z_2+Z_3=7\hat i+2\hat j"

Hence Z1,Z2 and Z3 are commutative under addition operation.


For distributive property, we must have...

"Z_i.(Z_j+Z_k)=Z_i.Z_j+Z_i.Z_k" for all i,j and k.

let's check...

"Z_1.(Z_2+Z_3)\stackrel{?}{=}Z_1.Z_2+Z_1.Z_3"

"LHS: (5\hat i-2\hat j).(7\hat i + 2 \hat j)=31\\
RHS: (15-6)+(20+2)=31"

Similarly we can show that...

"Z_2.(Z_3+Z_1)=Z_2.Z_3+Z_2.Z_1\\
Z_3.(Z_1+Z_2)=Z_3.Z_1+Z_3.Z_2"

Hence Z1, Z2 and Z3 follow distributive property.